A crate of mass 10kg is pulled up arough incline with an initial speed of 1.5m/s. the pulling force is 100N parallel to the incline which makes an angle of 30° with the horizontal. the coefficient of kinetic friction is 0.4 and the crate is pulled 5m. How much mechanical energy is lost due to friction?

1
Expert's answer
2021-12-07T20:11:33-0500

Energy loss is simply Kinetic friction multiply by the distance traveled. To find kinetic friction, take coefficient of friction multiply by normal force acting on the block.

"N = 10 \times 9.8 \times cos30 = 84.87 \\

N \times k = 84.87 \times 0.4 = 33.95 \\

Energy\; loss = 33.95 \times 5 = 169.74 \;J"